x^(2/3)-15x^(1/3)+26=0
Answer by palanisamy(496)

The given equation is
x^(2/3)-15x^(1/3)+26=0 ...(1)
Put x^(1/3) = y
Then, (1) becomes y^2-15y+26 = 0
(y-2)(y-13) = 0
y = 2 or y = 13
x^(1/3) = 2 or x^(1/3) = 13
x = 2^3 or x = 13^3
x = 8 or x = 2197